An English reading list ha 8 American novels and 6 English novels. A student must read 4 from the list and at least 2 must be American novels. In how many different ways can the four books be selected?
{This answer is 836. I'm obviously not doing it with the correct functions. It seems logical the C(8,2) are part of this problem. My answer is 1050 when I tried C(8,4)*C(6,4)
Any explanation for the correct way to do the work is greatly appreciated!

Answer by jim_thompson5910(35256)

There are C(8,2)*C(6,2) = 28*15 = 420 different ways to have exactly 2 American novels (and exactly two English novels)

There are C(8,3)*C(6,1) = 56*6 = 336 different ways to have exactly 3 American novels (and exactly one English novel)

and finally,

There are C(8,4)*C(6,0) = 70*1 = 70 different ways to have exactly 4 American novels (and no English novels)

Add these counts up to get


420+336+70 = 826

So there are 826 ways.
Please make sure that the answer you typed is not a typo (since the numbers look very close)